A corporate bond with $1000 maturity value carries a 7.5% coupon rate. It currently makes
interest payments semi-annually.
1
Expert's answer
2016-02-23T00:00:54-0500
RM1000 maturity value, 7.5% coupon rate, interest payments semi-annually.
If this 12-year bond currently sells for RM961.88, then the rate of return on this
bond is: Rate of return = (C + P1 - P0)/P0 = (1,000*0.075/2*24 + 961.88 - 1,000)/1,000 = 86.2%